# Multi-Agent Collaborative Data Selection for Efficient LLM Pretraining

### v1

@@ Abstract
Efficient data selection is crucial to accelerate the pretraining of large language models (LLMs). While various methods have been proposed to enhance data efficiency, limited research has addressed the inherent conflicts between these approaches to achieve optimal data selection for LLM pretraining. To tackle this problem, we propose a novel multi-agent collaborative data selection mechanism. In this framework, each data selection method serves as an independent agent, and an agent console is designed to dynamically integrate the information from all agents throughout the training process. We conduct extensive empirical studies to evaluate our multi-agent framework. The experimental results demonstrate that our approach significantly improves data efficiency, accelerates convergence in LLM training, and achieves an average performance gain up to $10.5\%$ across multiple language model benchmarks compared to the state-of-the-art methods. Code and checkpoints are publicly released at github.com/beccabai/multi-agent-data-selection.
@@ 1 Introduction
Efficient data selection is crucial for the pretraining of large language models (LLMs), as the quality of training data significantly impacts the statistical efficiency of the training procedure and the model performance (Brown, 2020; Du et al., 2022; Chowdhery et al., 2023). Recently, we have witnessed numerous approaches, such as filtering high-quality data (Xie et al., 2023b; Wettig et al., 2024), mixing data from multiple domains (Xie et al., 2023a; Liu et al., 2024), and selecting data that optimally boosts downstream task performance dynamically (Engstrom et al., 2024; Yu et al., 2024), which aim to improve data efficiency by prioritizing more informative training samples. However, these methods often operate independently or in isolated settings, limiting their potential when integrated into a collaborative framework. In this work, we want to explore how to effectively, flexibly, and robustly combine these advanced data selection techniques through the dynamic pretraining process, addressing the challenges of optimizing data efficiency for LLM pretraining at scale.
Nowadays, various heuristic methods have been proposed to provide measurements for the data samples used during LLM pre-training, aiming to optimize data efficiency by selecting or weighting the most informative training examples. However, we observe that integrating multiple data selection and mixing strategies presents significant challenges due to their inherent conflicts. For example, high-quality data identified by scoring functions may not align with data that strongly impact model performance as measured by influence functions (Engstrom et al., 2024); similar conflicts also exists between other methods — further details are enumerated in §2. These observations actually motivate us to launch a systematic discussion about how to effectively integrate these methods during the dynamic pretraining process that provides superior data efficiency for LLM pretraining.
On the other hand, effectively integrating these data selection methods into a single framework is much harder to implement than to ask for. In fact, one may have to explore an exponential space to find the optimal combination for different data sampling schemas. Such a heavy burden will be further amplified when we consider the dynamic adjustment during the training process introduced by state-of-the-art online data selection approaches. In fact, different from the offline methods that leverage fixed classifier-based scoring functions (Brown, 2020; Gao et al., 2020; Du et al., 2022; Chowdhery et al., 2023; Sachdeva et al., 2024; Wettig et al., 2024), domain weights (Brown, 2020; Team, 2024; Rae et al., 2021; Xie et al., 2023a; Liu et al., 2024), or down-sampled topics (Team, 2024; Chen et al., 2024), online methods use techniques like influence functions (Engstrom et al., 2024; Yu et al., 2024) to assess the model’s sensitivity to individual data points during the LLM pretraining process to enable dynamic selection of high-impact data at any optimization step. Such an online paradigm is computationally intensive by itself, which presents the essential challenge for an effective dynamic integration in our problem setting that demands a computationally efficient solution to preserve or even amplify the advantage of each data selection heuristic.
To address these challenges, we conduct a case study to identify the inherent conflicts for existing data selection methods and provide a multi-agent collaborative data selection framework to resolve this issue. Concretely, we make the following contributions:
Contribution 1: We conduct a case study in §2 over the SlimPajama dataset to present some interesting discoveries about the relationship among four popular data selection measurements used in LLM pretraining, i.e., data quality, topic diversity, data impact, and data domain, where the statistic suggests some interesting inherent conflicts among the existing methods.
Contribution 2: We propose a novel multi-agent collaborative data selection mechanism in §3. In this framework, each data selection method operates as an agent capable of providing scores to prioritize the training data samples. We also design an agent console to effectively integrate the scores from all agents, producing optimized data selection results. Furthermore, we implement a dynamic collaboration mechanism, where the contribution of each agent can be adjusted dynamically throughout the LLM training process. This approach enables more flexible and adaptive data selection, improving overall data efficiency during LLM pretraining.
@@
Figure 1: Statistics of the SlimPajama dataset. This figure shows the distribution of overall 670B tokens data across four dimensions: quality, domain, topic diversity, and impact on the pretrained model at the $1500$th step. Each bar represents a subset defined by a specific quality interval and domain. This visualization reveals the conflict among diversity, quality, and model influence.
Contribution 3: To evaluate our multi-agent collaborative data selection method, we conducted extensive experiments to show that: (1) In end-to-end experiments, our approach significantly improves data efficiency, leading to faster convergence in LLM training and achieving an average improvement up to $10.5\%$ across various language model benchmarks compared to baseline methods (§4.1); and (2) Ablation studies confirm that the design and implementation of key components in our multi-agent framework are crucial for attaining this advanced performance (§4.2). These findings highlight the effectiveness of our method to optimize data efficiency for LLM training.
@@ 2 Case Study - Inherent Conflicts in Data Selection
We believe this inherent conflict illustrates that a naive ensemble of these mechanisms may lead to poor performance in terms of data efficiency for LLM pretraining, which motivates the design and implementation of our multi-agent collaborative framework in §3.
3 Multi-agent Collaborative Data Selection
In this section, we present the formalization of the data selection problem in §3.1, outline the overall framework of our methods in §3.2, and detail the agent initialization and update in §3.3, along with the collaborative mechanism in §3.4.
@@ 3.1 Problem Formulation
$\mathcal{D}_{k}^{*}$ $\coloneqq\operatorname*{arg\,min}_{\mathcal{D}_{k}\subset\mathcal{D},|\mathcal{D}_{k}|=k}\mathcal{L}(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{D}_{\text{eval}}),$ (1)
$\text{where }\mathcal{L}(\mathcal{D}_{k}\mid$ $\mathcal{M},\mathcal{D}_{\text{eval}})\coloneqq\mathbb{E}_{x\sim\mathcal{D}_{\text{eval}}}\left[\ell(x;\mathcal{O}(\mathcal{M},\mathcal{D}_{k}))\right],$
@@
where $\mathcal{D}_{\text{eval}}$ denotes the downstream tasks and $\ell(\cdot)$ denotes the loss function on the data sample $x$. Minimizing this objective directly is computationally challenging. Given that the real downstream tasks are unknown during model training, prior works have approximately optimized this problem by minimizing the loss on selected reference tasks $D_{\text{ref}}$ (e.g. LAMBADA Paperno et al. (2016), SQuAD Rajpurkar (2016) and Jeopardy Tunguz (2019) in Engstrom et al. (2024)). Specifically, they train proxy models to compute one-step training loss (Yu et al., 2024) or influence functions (Engstrom et al., 2024) on the reference tasks to approximate the true loss. However, this approach heavily depends on the selection of the reference tasks, while the chosen reference tasks may not be fully representative of all potential downstream tasks.
@@
Figure 2: Illustration of multi-agent collaborative framework. Multi-agent collaborative framework for pretraining data selection that integrates multiple perspectives by combining offline priors and online model-derived preferences.
@@
To avoid this obstacle, we do not directly minimize the loss on the reference tasks. Instead, we view this loss as a reward signal that guides the update of predefined data selection methods. Concretely, we define a reward function $R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{D}_{\text{ref}})$, where the reward is based on the performance gain of current model $\mathcal{M}$ trained on the subset $\mathcal{D}_{k}$ and evaluated on the reference tasks $\mathcal{D}_{\text{ref}}$. Then our optimization goal becomes maximizing this reward over time, as:
$\mathcal{D}_{k}^{*}$ $=\operatorname*{arg\,max}_{\mathcal{D}_{k}\subset\mathcal{D},|\mathcal{D}_{k}|=k}\mathbb{E}\left[R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{D}_{\text{ref}})\right],$ (2)
where $R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{D}_{\text{ref}})\coloneqq\mathbb{E}_{x\sim\mathcal{D}_{\text{ref}}}\left[-\ell(x;\mathcal{O}(\mathcal{M},\mathcal{D}_{k}))\right].$
3.2 Multi-agent Collaborative Data Selection Framework
@@
In order to solve the optimization problem in Equation 2, we develop a framework illustrated in Figure 2. This framework consists of two primary stages: the offline labeling stage and the online update stage. Before the training process, some initial information (i.e., the initialized measurements in some heuristic) is computed for the entire pretraining corpus, and this information is stored separately in each agent’s memory (formally defined below). During the training process, the current model (i.e., LLM to be trained) is used to update the agents’ memory and their collaboration mechanism based on rewards computed on the current model. An agent console is responsible for aggregating the opinions of each agent and making the final data selection decision. Formally, we define the agent in Definition 1 and the agent console in Definition 2.
Definition 1 (Agent).
@@
An agent $\mathcal{A}$ is a data selection method defined by a specific attribute (e.g., quality, domain, or topic) with memory $\mathcal{H}_{\mathcal{A}}$ that stores labels for each data point and their associated scores. During training, the agent samples data $\mathcal{D}_{\mathcal{A}}$, calls the current model to compute the reward $\mathcal{R}(x_{i})$ for each sample $x_{i}\in\mathcal{D}_{\mathcal{A}}$, and updates its internal weights $\mathbf{w_{\mathcal{A}}}$ that determine the score for each data point based on the rewards calculated by this particular agent. The agent then assigns a score $\mathcal{S}_{\mathcal{A}}$ to each data point based on its memory and the updated weights. One agent’s objective is to maximize its reward $R(\mathcal{D}_{\mathcal{A}}\mid\mathcal{M},\mathcal{D}_{\text{ref}})$ by updating this agent’s internal weights and increasing the score of higher-reward data points.
Definition 2 (Agent Console).
@@
The agent console is a coordination mechanism that consolidates scores $\mathcal{S}_{\mathcal{A}}(x_{i})$ from multiple agents $\{\mathcal{A}_{1},\dots,\mathcal{A}_{n}\}$ to calculate a final score $S(x_{i})$ for each data point $x_{i}$. The console adjusts the collaborative weights $\theta_{\mathcal{A}}$ for each agent based on their respective aggregate rewards $\mathcal{R}_{\mathcal{A}}$, balancing their contributions during training. In cases where there are conflicts in the decisions made by agents, the console resolves these by adjusting the weights $\theta_{\mathcal{A}}$ to prioritize the agents that have a greater positive impact on the model’s performance, ensuring an effective data selection process.

### v3

# Efficient Pretraining Data Selection for Language Models via Multi-Actor Collaboration

@@ Abstract
Efficient data selection is crucial to accelerate the pretraining of language model (LMs). While various methods have been proposed to enhance data efficiency, limited research has addressed the inherent conflicts between these approaches to achieve optimal data selection for LM pretraining. To tackle this problem, we propose a multi-actor collaborative data selection mechanism: each data selection method independently prioritizes data based on its criterion and updates its prioritization rules using the current state of the model, functioning as an independent actor for data selection; and a console is designed to adjust the impacts of different actors at various stages and dynamically integrate information from all actors throughout the LM pretraining process. We conduct extensive empirical studies to evaluate our multi-actor framework. The experimental results demonstrate that our approach significantly improves data efficiency, accelerates convergence in LM pretraining, and achieves an average relative performance gain up to $10.5\%$ across multiple language model benchmarks compared to the state-of-the-art methods. Code and checkpoints are publicly released at https://github.com/Relaxed-System-Lab/multi-actor-data-selection.
Efficient Pretraining Data Selection for Language Models via Multi-Actor Collaboration
Tianyi Bai1,2, Ling Yang3, Zhen Hao Wong3, Fupeng Sun5, Xinlin Zhuang2, Jiahui Peng2, Chi Zhang2, Lijun Wu 2, Jiantao Qiu2thanks: Corresponding authors.thanks: Project lead., Wentao Zhang2,3,4*, Binhang Yuan1*, Conghui He2*, 1Hong Kong University of Science and Technology, 2Shanghai Artificial Intelligence Laborator, 3Peking University, 4Zhongguancun Academy, 5Imperial College London Correspondence: wentao.zhang@pku.edu.cn, biyuan@ust.hk, {heconghui, qiujiantao}@pjlab.org.cn
@@ 1 Introduction
Efficient data selection is crucial for the pretraining of language model (LMs), as the quality of training data significantly impacts the statistical efficiency of the training procedure and the model performance (Brown et al., 2020; Du et al., 2022; Chowdhery et al., 2023). Recently, we have witnessed numerous approaches, such as filtering high-quality data (Xie et al., 2023b; Wettig et al., 2024), mixing data from multiple domains (Xie et al., 2023a; Liu et al., 2024), and selecting data that optimally boosts downstream task performance dynamically (Engstrom et al., 2024; Yu et al., 2024), which aim to improve data efficiency by prioritizing more informative training samples. However, these methods often operate independently or in isolated settings, limiting their potential when integrated into a collaborative framework. In this work, we want to explore how to effectively, flexibly, and robustly combine these advanced data selection techniques through the dynamic pretraining process, addressing the challenges of optimizing data efficiency for LM pretraining at scale.
@@
Figure 1: This figure shows the 627B token distribution by quality, domain, topic diversity, and model impact at step 1500. Each bar represents a subset, highlighting trade-offs between diversity, quality, and influence.
Nowadays, various heuristic methods have been proposed to provide measurements for the data samples used during LM pre-training, aiming to optimize data efficiency by selecting or weighting the most informative training examples. However, we observe that integrating multiple data selection and mixing strategies presents significant challenges due to their inherent conflicts. For example, high-quality data identified by scoring functions may not align with data that strongly impact model performance as measured by influence functions (Engstrom et al., 2024); similar conflicts also exists between other methods — further details are enumerated in §2. These observations actually motivate us to launch a systematic discussion about how to effectively integrate these methods during the dynamic pretraining process that provides superior data efficiency for LM pretraining.
Effectively integrating multiple data selection methods into a unified framework presents significant challenges. While each method may individually offer benefits, combining them requires navigating an exponential search space of possible configurations, which quickly becomes infeasible at scale. This challenge is further exacerbated in online data selection settings, where decisions must be made dynamically during training. Unlike offline approaches that rely on static, precomputed heuristics—such as classifier-based scoring, domain weighting, or topic sampling (Brown et al., 2020; Team, 2024; Xie et al., 2023a)—online methods aim to adaptively select data based on the model’s evolving learning state. For example, MATES (Yu et al., 2024) continuously probes the pretraining model to estimate the influence of individual data points and trains a lightweight data influence model to predict which samples will be most beneficial at each stage of training. Similarly, DSDM (Engstrom et al., 2024) formulates data selection as an optimization problem, selecting data subsets that maximize performance on downstream tasks without relying on predefined quality metrics. Despite their effectiveness, these methods incur substantial computational overhead, as they require labeling or evaluating the entire dataset at each stage of training—an approach that is impractical for large-scale pretraining. Consequently, the core challenge is to develop an online integration framework that preserves the adaptivity and performance benefits of these techniques while remaining computationally efficient and scalable.
To address these challenges, we conduct a case study to identify the inherent conflicts for existing data selection methods and provide a multi-actor collaborative data selection framework to resolve this issue. Our multi-actor framework is inspired by the classical definition of intelligence outlined by (Russell and Norvig, 2016), where an actor is defined as an entity that perceives a state and maps the observed state to actions. In our approach, data selection is achieved by the collaboration of these actors. Our contributions can be summarized as:
Contribution 1: Our case study on the SlimPajama reveals conflicts among four data selection metrics in LM pretraining. Despite these conflicts, prior studies (Wettig et al., 2024; Xie et al., 2023a; Yu et al., 2024) show that even a single metric can effectively guide training, highlighting the need for better integration of these approaches.
Contribution 2: We propose a novel multi-actor collaborative data selection framework (§3), where each method acts as an independent scorer for training data. An actor console integrates these scores to optimize selection, and a dynamic collaboration mechanism adjusts actor contributions throughout training, enhancing flexibility and data efficiency.
Contribution 3: Extensive experiments demonstrate that: (1) our method significantly improves data efficiency, accelerating LM convergence and achieving up to a 10.5% performance boost over baselines (§4.1); and (2) ablation studies confirm that key components of our framework are essential for these gains (§4.2).
@@ 2 Case Study - Inherent Conflicts in Data Selection
We believe this inherent conflict illustrates that a naive ensemble of these mechanisms may lead to poor performance in terms of data efficiency for LM pretraining, which motivates the design and implementation of our multi-actor collaborative framework in §3.
3 Collaborative Data Selection
In this section, we present the formalization of the data selection problem in §3.1, outline the overall framework of our methods in §3.2, and detail the actor initialization and update in §3.3, along with the collaborative mechanism in §3.4.
@@ 3.1 Problem Formulation
$\mathcal{D}_{k}^{*}$ $\coloneqq\operatornamewithlimits{arg\,min}_{\mathcal{D}_{k}\subset\mathcal{D},|\mathcal{D}_{k}|=k}\mathcal{L}(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{T}_{\text{eval}}),$ (1)
where
$\mathcal{L}(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{T}_{\text{eval}})\coloneqq\mathbb{E}_{x\sim\mathcal{T}_{\text{eval}}}\left[\ell(x;\mathcal{O}(\mathcal{M},\mathcal{D}_{k}))\right]$
@@
denotes the expected loss (e.g., cross-entropy loss) for model $\mathcal{M}$ on downstream task $\mathcal{T}_{\text{eval}}$. Minimizing this objective directly is computationally challenging. Given that the real downstream tasks are unknown during model training, prior works have approximately optimized this problem by minimizing the loss on selected reference tasks $D_{\text{ref}}$ (e.g. LAMBADA Paperno et al. (2016), SQuAD Rajpurkar (2016) and Jeopardy Tunguz (2019) in Engstrom et al. (2024)). Specifically, they train proxy models to compute one-step training loss (Yu et al., 2024) or influence functions (Engstrom et al., 2024) on the reference tasks to approximate the true loss. However, this approach heavily depends on the selection of the reference tasks, while the chosen reference tasks may not be fully representative of all potential downstream tasks.
@@
Figure 2: Illustration of multi-actor collaborative framework. Multi-actor collaborative framework for pretraining data selection that integrates multiple perspectives by combining offline priors and online model-derived preferences.
@@
To avoid this obstacle, we do not directly minimize the loss on the reference tasks. Instead, we view this loss as a reward signal that guides the update of predefined data selection methods. Concretely, we define a reward function $R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{T}_{\text{ref}})$, where the reward is based on the performance gain of current model $\mathcal{M}$ trained on the subset $\mathcal{D}_{k}$ and evaluated on the reference tasks $\mathcal{T}_{\text{ref}}$. Then our optimization goal becomes maximizing this reward over time, as:
$\mathcal{D}_{k}^{*}=\operatornamewithlimits{arg\,max}_{\mathcal{D}_{k}\subset\mathcal{D},|\mathcal{D}_{k}|=k}\mathbb{E}\left[R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{T}_{\text{ref}})\right]$ (2)
where
$R(\mathcal{D}_{k}|\mathcal{M},\mathcal{T}_{\text{ref}})\coloneqq\mathbb{E}_{x\sim\mathcal{T}_{\text{ref}}}\left[-\ell(x;\mathcal{O}(\mathcal{M},\mathcal{D}_{k}))\right].$
We will rewrite $R(\mathcal{D}_{k}\mid\mathcal{M},\mathcal{T}_{\text{ref}})$ as $R(\mathcal{D}_{k})$ if there is no ambiguity. In practice, $R(\mathcal{D}_{k})$ can be approximated by the weighted average of influence functions Engstrom et al. (2024); Yu et al. (2024), which is defined by
$r(x_{i})=-\nabla_{\mathcal{M}}\mathcal{L}(\mathcal{T}_{\text{ref}}\mid\mathcal{M})^{\top}H^{-1}_{\mathcal{M}}\nabla_{\mathcal{M}}\mathcal{L}(x_{i}\mid{\mathcal{M}}),$ (3)
where $H_{\mathcal{M}}=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}_{\mathcal{M}}\mathcal{L}_{\mathcal{M}}(x_{i}\mid{\mathcal{M}})$ is the Hessian and its positive definite. Details of calculating influence functions for pretraining data point can be found in §A.9.
3.2 Multi-Actor Data Selection Framework
@@
In order to solve the optimization problem in Equation 2, we develop a framework illustrated in Figure 2. This framework consists of two primary stages: the offline labeling stage and the online update stage. Before the training process, some initial information (i.e., the initialized measurements in some heuristic) is computed for the entire pretraining corpus, and this information is stored separately in each actor’s memory (formally defined below). During the training process, the current model (i.e., LM to be trained) is used to update the actors’ memory and their collaboration mechanism based on rewards computed on the current model. An actor console is responsible for aggregating the opinions of each actor and making the final data selection decision. Formally, we define the actor in Definition 1 and the actor console in Definition 2. Detailed formulation is in Appendix A.4.
Definition 1 (Actor).
@@
An actor $\mathcal{A}$ is a data selection method defined by a specific attribute (e.g., quality, domain, or topic) with memory $\mathcal{H}_{\mathcal{A}}$ that stores labels for each data point and their associated scores. During training, the actor takes several actions: (1) Sample data $\mathcal{D}_{\mathcal{A}}$ according to predefined sampling distribution, (2) Call the current model to compute the reward $\mathcal{R}(\mathcal{D}_{\mathcal{A}})$, (3) Get feedbacks from current model state, and (4) Update the internal weights $\mathbf{w_{\mathcal{A}}}$ in its memory. Then it assigns a score $\mathcal{S}_{\mathcal{A}}$ to each data point based on its updated memory, prioritizing the good data according to the updated weights. One actor’s objective is to maximize its reward by updating this actor’s internal weights and increasing the score of higher-reward data points.
Definition 2 (Actor Console).
@@
The actor console is in charge of coordinate opinions from different actor to make final decision of selecting dataset for next training stage. Specifically, it consolidates scores $\mathcal{S}_{\mathcal{A}}(x_{i})$ from multiple actors $\{\mathcal{A}_{1},\dots,\mathcal{A}_{n}\}$ to calculate a final score $S(x_{i})$ for each data point $x_{i}$, and select the final dataset. The console adjusts the collaborative weights $\theta_{\mathcal{A}}$ for each actor based on their respective aggregate rewards $\mathcal{R}(\mathcal{D}_{\mathcal{A}})$, balancing their contributions during training. In cases where there are conflicts in the decisions made by actors, the console resolves these by adjusting the weights $\theta_{\mathcal{A}}$ to prioritize the actors that have a greater positive impact on the model’s performance, ensuring an effective data selection process.